# Non-stabilizerness versus entanglement in matrix product states

### v1

@@ I Introduction
The past two decades have witnessed a surge in leveraging quantum information tools to tackle the long-standing challenge of the many-body problem [1]. This synergy has opened avenues for novel theoretical frameworks and powerful computational techniques to describe and simulate complex quantum systems [2]. Entanglement is a key resource for performing several quantum information protocols [3]. Building on its early understanding in few-body systems, over the last two decades, its importance for the many-body problem has been uncovered [4, 5, 6]. It has found applications over a wide range of phenomena, from real-time dynamics [7], to topological order [8, 9] and classification of states [10, 11]. Moreover, a particularly useful perspective has pivoted on the role of bipartite entanglement in characterizing quantum wave functions, leading to the formulation of several classes of tensor network (TN) states [12, 13, 14, 15].
@@
However, it is also well known that entanglement is not the only resource required for specific tasks. In particular, in the context of quantum computing, an equally pivotal resource is non-stabilizerness - also known as magic [16, 17]. Much like entanglement, nonstabilizerness has been quantified within the framework of resource theory using measures of nonstabilizerness. Several measures of nonstabilizerness have been proposed in quantum information theory, with most of them relying on the notion of quasiprobability distributions [18, 19, 20, 21, 22]. However, most of these quantifiers are difficult to evaluate even numerically (see, e.g., Refs. [23, 24, 25, 26, 27, 28]). This computational intractability has hindered the task of quantifying nonstabilizerness beyond a few qubits, posing a major challenge in the field. To address this challenge, computable and practical measures of nonstabilizerness have been introduced recently: the Bell magic [29], the stabilizer nullity [30] and the stabilizer Renyi entropies (SREs) [31]. Notable progress has been made in their experimental measurements [32, 33, 34, 35, 36].
@@
Very much like when entanglement entered the many body realm, it is presently not very clear how much magic relates to physical phenomena, and how magic and entanglement are intertwined at the many-particle level. A very promising framework where the above question might be answered is indeed provided by TNs. The control parameter describing their descriptive power, often referred to as bond dimension, is directly connected with the entanglement. Most relevantly, TNs provide presently the only theoretical framework where magic can be computed at large scale, utilizing both exact [37, 38] and stochastic methods[39, 40, 41, 42].
@@
This work delves into the connection between nonstabilizerness and the bond dimension of matrix product states (MPSs) [43]. We pursue a two-pronged approach. First, we investigate the broader question of how these two quantum resources, crucial for realizing states intractable to classical computers, are interrelated in many-body systems. Both nonstabilizerness and entanglement are known to be essential for such simulations. Second, we focus on the convergence properties of magic within the context of MPS simulations. Essentially, for a given variational simulation with a fixed bond dimension $\chi$, we explore how quickly these approximations converge to the true values. Particularly, we address the magic-bond dimension relationship in two distinct, yet equally intriguing, many-body scenarios: full state magic and mutual magic of ground states (GSs). These scenarios offer diverse perspectives on the interplay between magic and entanglement. In the first case, we compare a global property (magic) with correlations between arbitrary sub-system parts (entanglement). Conversely, mutual magic examines shared resources between partitions, potentially relevant to field theory due to the cancellation of all boundary terms inherent to the chosen lattice regularization.
@@ II.1 Resource theory for magic
$\mathcal{P}_{N}=\left\{e^{i\theta\frac{\pi}{2}}\sigma_{1}\otimes\cdots\sigma_{j_{N}}|\theta,j_{k}=0,1,2,3\right\}.$ (1)
@@
In this context, it is important to quantify the non-Clifford resources needed to prepare a given state. The amount of non-stabilizerness, or magic, of any state is measured using magic monotones. The properties required for a good monotone $\mathcal{M}$ of nonstabilizerness are: (i) $\mathcal{M}(|\psi\rangle)=0$ iff $|\psi\rangle$ is a stabilizer state (ii) nonincreasing under Clifford operations: $\mathcal{M}(\Gamma|\psi\rangle)<\mathcal{M}(|\psi\rangle)$, and (iii) $\mathcal{M}(|\psi\rangle\otimes|\phi\rangle)=\mathcal{M}(|\psi\rangle)+\mathcal{M}(|\phi\rangle)$ (in general, sub-additivity is sufficient).
Previous studies primarily explored magic measures in small or weakly correlated systems, hindering our grasp of nonstabilizerness in entangled, many-body systems. This limitation stems from the exponential growth of stabilizer states and their increasingly complex geometries with increasing system size. Consequently, calculating or numerically analyzing nonstabilizerness measures for large states becomes tremendously challenging. Only very recently this has been made possible by the developments of analytical and numerical tools to compute the SREs at large scales [37, 38, 39, 40, 41, 42, 50, 51, 52, 53].
From a quantum information perspective, understanding and measuring nonstabilizerness in MPS is crucial because it’s a potential resource for achieving quantum advantage. Recent studies suggest that the computational power of a state isn’t solely determined by its overall nonstabilizerness (magic density). Other characteristics, including subleading terms, nonlocal components, and potentially topological aspects, also play significant roles.
@@ II.2 Stabilizer Rényi entropies
with $d$ is the local dimension of the Hilbert space of $N$ qudits and $\mathcal{P}_{N}$ is the generalized Pauli group of $N$ qudits. The SREs have the following properties: [31] (i) faithfulness: $M_{n}(\rho)=0$ iff $\rho\in\text{STAB}$, (ii) stability under Clifford unitaries $C\in\mathcal{C}_{N}$: $M_{n}(C\rho C^{\dagger})=M_{n}(\rho)$ , and (iii) additivity: $M_{n}(\rho_{A}\otimes\rho_{B})=M_{n}(\rho_{A})+M_{n}(\rho_{B})$. The SREs are thus a good magic measure in the point of view of resource theory, where the free states are defined as the stabilizer states while the free operations are the Clifford unitaries. This definition is a straightforward generalization to general local dimension $d$ from the one given in Ref. [31]. For $d>2$, the Pauli operators are no longer Hermitian, and thus the expectation values can be complex. For this reason, we take the absolute values of the expectation values $|\text{Tr}\left(\rho P\right)|^{2n}$. Eq.(3) can be seen as the Rényi-n entropy of the classical probability distribution:
@@ II.4.1 Perfect Pauli sampling for MPS
This method is based on the fact that the SREs can be evaluated as
@@
where $p(\alpha_{j}|\alpha_{1}\alpha_{2}\cdots\alpha_{j-1})$ is the conditional probability to get a Pauli matrix $\sigma^{\alpha_{j}}$ at site $j$, given the partial configuration $\sigma_{1}^{\alpha_{1}}\cdots\sigma_{j-1}^{\alpha_{j-1}}$ and marginalising over the remaining lattice sites. By iterating over this chain relation, each conditional probability can be efficiently computed for a MPS. In particular, after a complete sweep along the entire system, and a total computational cost $O(N\chi^{3})$, we generate strings of Pauli operators $P_{\bm{\alpha}}$ associated exactly with the correct probability $\Xi_{P_{\bm{\alpha}}}$ [39, 40]. Then $M_{n}$ (or $M_{1}$) can be estimated using Eq.s (11,12) averaging over a finite number of samples $\mathcal{S}$; the accuracy of the method of course increases with the number of samples. Indeed, the difference between the average value of $M_{n}$ obtained by sampling the distribution $\Xi_{P}$ and the actual value $M_{n}$ is of the order of $O(1/\sqrt{S})$. In addition, in the typical cases where the SREs is extensive, the number of samples required to estimate the SREs within a given error scales polynomially with $N$ for $n=1$, and it is exponential in $N$ for $n\neq 1$. [39, 40, 41]. Finally, it’s worth mentioning that perfect Pauli sampling for MPS can be enhanced by employing a clever biasing strategy, enabling direct access to the stabilizer nullity [57], which can be seen as well as the $n\to\infty$ limits of the SREs [38].
@@ III SRE in spin-1 XXZ chain
where $S^{\alpha}$’s, $\alpha=x,y,z$, are the spin-1 operators, $J_{z}$ is the easy-axis anisotropy, and $D$ is the single-ion anisotropy. The model has a global $U(1)$ symmetry corresponding to the conservation of total magnetization $\sum_{i}S^{z}_{i}$, and here we consider the ground states at the sector of zero total magnetization. The model displays a rich phase diagram, making it a good playground to explore magic in phase transitions [58].
@@ III.1 Scaling of full-state magic
Moreover, in Fig. 3 and Fig. 4, we demonstrate that $m_{1}$ scales as $\frac{1}{\chi^{2}}$ at critical points. Fig. 3 shows the linear dependence of $m_{1}$ from $\frac{1}{\chi^{2}}$ for different system sizes ($N\in\{128,256,512,1024\}$), while Fig. 4 provides a detailed view of the largest system size. The dashed lines in these figures represent the fits performed with the functional form $m_{1}(1/\chi^{2})=m_{0}+c~{}1/\chi^{2}$. Note that the slope $c$ is size independent in the Large D-XY and Haldane-Neel transitions, while it changes non-monotonically with $N$ in the Haldane-Large D case. Regarding the intercepts, we expect them to be size invariant since the SRE is generally linear with size $N$ and what we’re computing is its density, namely $m_{1}=M_{1}/N$. Indeed, this expectation holds true for the Large D-XY and Haldane-Large D transitions but not for the Haldane-Néel one, which appears to exhibit a more pronounced finite-size effect.
@@
Moreover, in Fig. 5, we show the behavior of $m_{2}$ as a function of bond dimension $\chi$ obtained from Pauli-MPS with bond dimension $\chi_{P}=2*\chi$. In Fig. 5a we show $m_{2}$ vs $\chi$ in the three distinct transitions: Haldane-Neel, Haldane-Large F and Large D-XY. We observed that the convergence of the $m_{2}$ occurs at small bond dimension. In Fig 5b, we show that $m_{2}$ scales as $\frac{1}{\chi^{2}}$ at the critical points for system size $N=64$. In the three different panels, the dashed lines represent the fits performed with the functional form $m_{2}(1/\chi^{2})=m_{0}+c~{}1/\chi^{2}$.
@@ III.2 Mutual information and long-range magic scaling
As concerns long-range magic, the numeric results are challenging to interpret due to statistical error. In Fig. 8a and Fig. 9a, we display the scaling with the system size $N$ for case $BC$ and $AC$, respectively. Comparing with the behavior of mutual information, long-range magic appears to increase less rapidly for two connected partitions, while remaining constant for disconnected partitions. Turning to the scaling with the bond dimension $\chi$, depicted in Fig. 8b, Fig. 9b, we notice parallels with the behavior of mutual information. However, unlike for the full-state magic case, we are not able to predict a functional form for the correction in the bond dimension. Specifically, to estimate the $\chi$-dependence, in the previous section we focused on $\chi\geq 5$. Yet, at these values of $\chi$, it is difficult to discern the scaling, as the fluctuations we aim to fit are of the same order as the error bars with a sample size of $N_{S}=10^{6}$. This is because our estimation of long-range magic comprises a sum of two terms, each associated with its own statistical error, and the $W$ term defined in Eq. (16) is more challenging to estimate. Consequently, we cannot predict the precise scaling of long-range magic in $\chi$, but we consistently observe rapid saturation.
@@ Acknowledgments
We thank A. Hamma, T. Haug, G. Lami and L. Piroli for useful discussions and feedback on the manuscript, and M. C. Bañuls, R. Fazio and G. Fux for collaboration on related topics. P.S.T. acknowledges support from the Simons Foundation through Award 284558FY19 to the ICTP. M. D. was partly supported by the MIUR Programme FARE (MEPH), by the EU-Flagship programme Pasquans2, by the PNRR MUR project PE0000023-NQSTI, and by the PRIN programme (project CoQuS). M.C. was partially supported by the PRIN 2022 (2022R35ZBF) - PE2 - “ManyQLowD”. MD would like to thank the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) for their support. MD and ET were partly supported by QUANTERA DYNAMITE PCI2022-132919,
Note added: while completing this work, we became aware of a related study in the context of random matrix product states by Lami, Haug and De Nardis, that will appear on the same arxiv posting.
@@ Appendix A Autocorrelations
To further illustrate the origin of the above results, in Fig. 12, we show the behavior of the normalized autocorrelation function of the mutual information $I$ for case $BC$. Figs. 12a, 12b and 12c correspond, respectively, to the Large D-XY, Haldane-Large D and Haldane-Néel transitions. Counterintuitively, we observe that despite the mutual information increasing with size (see Fig. 6), the autocorrelation of the Pauli-Markov chain remains constant.

### v2

@@ I Introduction
The past two decades have witnessed a surge in leveraging quantum information tools to tackle the long-standing challenge of the many-body problem [1]. This synergy has opened avenues for novel theoretical frameworks and powerful computational techniques to describe and simulate complex quantum systems [2]. Entanglement, a key resource for performing several quantum information protocols [3], has played a pivotal role in this context [4, 5, 6]. Entanglement has been discussed in the characterization of a wide range of phenomena, from real-time dynamics [7], to topological order [8, 9] and classification of states [10, 11]. Moreover, a particularly useful perspective has pivoted on the role of bipartite entanglement in characterizing quantum wave functions, leading to the formulation of several classes of tensor network (TN) states [12, 13, 14, 15].
@@
However, it is also well known that entanglement is not the only resource required for specific quantum information tasks. In particular, in the context of quantum computing, an equally essential resource is non-stabilizerness - also known as magic [16, 17]. Much like entanglement, nonstabilizerness has been quantified within the framework of resource theory using measures of nonstabilizerness. Several measures of nonstabilizerness have been proposed in quantum information theory, with most of them relying on the notion of quasiprobability distributions [18, 19, 20, 21, 22]. However, in the context of many-body systems, most of these quantifiers are difficult to evaluate even numerically (see, e.g., Refs. [23, 24, 25, 26, 27, 28]), as they require in principle very complex minimization procedures. This computational intractability has hindered the task of quantifying nonstabilizerness beyond a few qubits, posing a major challenge in the field. To address this challenge, computable and practical measures of nonstabilizerness have been introduced recently: Bell magic [29], stabilizer nullity [30] and stabilizer Renyi entropies (SREs) [31]. Notable progress has been made in their experimental measurements [32, 33, 34].
@@
Very much like when entanglement entered the many body realm, it is presently not very clear how (and if) magic relates to physical phenomena, and how magic and entanglement are intertwined at the many-particle level. A very promising framework where the above question might be answered is provided by TN states. The control parameter describing their descriptive power, often referred to as bond dimension, is directly connected with the entanglement. Most relevantly, TN states provide presently the only theoretical framework where magic can be computed at large scale, utilizing both exact [35, 36] and stochastic methods[37, 38, 39, 40].
@@
Here, we study the connection between nonstabilizerness and the bond dimension of matrix product states (MPSs) [41]. We pursue a two-pronged approach. First, we investigate the broader question of how these two quantum resources, crucial for realizing states intractable to classical computers, are interrelated in many-body systems. Both nonstabilizerness and entanglement are known to be essential for such simulations. Second, we focus on the convergence properties of magic within the context of MPS simulations. Essentially, for a given variational simulation with a fixed bond dimension $\chi$, we explore how quickly these approximations converge to the true values.
We address the magic-bond dimension relationship in two distinct, yet equally intriguing, many-body scenarios: full state magic and mutual magic of ground states (GSs). These scenarios offer diverse perspectives on the interplay between magic and entanglement. In the first case, we compare a global property (magic) with correlations between arbitrary sub-system parts (entanglement). Conversely, mutual magic examines shared resources between partitions, potentially relevant to field theory due to the cancellation of all boundary terms inherent to the chosen lattice regularization.
@@ II.1 Resource theory for magic
$\mathcal{P}_{N}=\left\{e^{i\theta\frac{\pi}{2}}\sigma_{j_{1}}\otimes\cdots\sigma_{j_{N}}|\theta,j_{k}=0,1,2,3\right\}.$ (1)
@@
In this context, it is important to quantify the non-Clifford resources needed to prepare a given state. The amount of non-stabilizerness, or magic, of any state is measured using magic monotones. The properties required for a good monotone $\mathcal{M}$ of nonstabilizerness are: (i) $\mathcal{M}(|\psi\rangle)=0$ iff $|\psi\rangle$ is a stabilizer state (ii) nonincreasing under Clifford operations: $\mathcal{M}(\Gamma|\psi\rangle)\leq\mathcal{M}(|\psi\rangle)$, and (iii) $\mathcal{M}(|\psi\rangle\otimes|\phi\rangle)=\mathcal{M}(|\psi\rangle)+\mathcal{M}(|\phi\rangle)$ (in general, sub-additivity is sufficient).
Previous studies primarily explored magic measures in small or weakly correlated systems, hindering our grasp of nonstabilizerness in entangled, many-body systems. This limitation stems from the exponential growth of stabilizer states and their increasingly complex geometries with increasing system size. As a consequence, calculating or numerically analyzing nonstabilizerness measures for large states becomes tremendously challenging. Only very recently this has been made possible by the developments of analytical and numerical tools to compute the SREs at large scales [35, 36, 37, 38, 39, 40, 48, 49, 50, 51].
From a quantum information perspective, understanding and measuring nonstabilizerness in MPS is crucial because it’s a potential resource for achieving quantum advantage [45, 49, 52]. Recent studies suggest that the computational power of a state isn’t solely determined by its overall nonstabilizerness (magic density). Other characteristics, including subleading terms [39], nonlocal components [49], and potentially topological aspects [53], also play significant roles.
@@ II.2 Stabilizer Rényi entropies
with $d$ is the local dimension of the Hilbert space of $N$ qudits and $\mathcal{P}_{N}$ is the generalized Pauli group of $N$ qudits. The SREs have the following properties: [31] (i) faithfulness: $M_{n}(\rho)=0$ iff $\rho\in\text{STAB}$, (ii) stability under Clifford unitaries $C\in\mathcal{C}_{N}$: $M_{n}(C\rho C^{\dagger})=M_{n}(\rho)$ , and (iii) additivity: $M_{n}(\rho_{A}\otimes\rho_{B})=M_{n}(\rho_{A})+M_{n}(\rho_{B})$. The SREs are thus a good magic measure from the point of view of resource theory, where the free states are defined as the stabilizer states while the free operations are the Clifford unitaries. This definition is a straightforward generalization to general local dimension $d$ from the one given in Ref. [31]. For $d>2$, the Pauli operators are no longer Hermitian, and thus the expectation values can be complex. For this reason, we take the absolute values of the expectation values $|\text{Tr}\left(\rho P\right)|^{2n}$. Eq.(3) can be seen as the Rényi-n entropy of the classical probability distribution:
@@ II.4.1 Perfect Pauli sampling for MPS
This method is based on Tensor Networks perfect sampling algorithms [57, 58] and on the fact that the SREs can be evaluated as
@@
where $p(\alpha_{j}|\alpha_{1}\alpha_{2}\cdots\alpha_{j-1})$ is the conditional probability to get a Pauli matrix $\sigma^{\alpha_{j}}$ at site $j$, given the partial configuration $\sigma_{1}^{\alpha_{1}}\cdots\sigma_{j-1}^{\alpha_{j-1}}$ and marginalising over the remaining lattice sites. By iterating over this chain relation, each conditional probability can be efficiently computed for a MPS. In particular, after a complete sweep along the entire system, and a total computational cost $O(N\chi^{3})$, we generate strings of Pauli operators $P_{\bm{\alpha}}$ associated exactly with the correct probability $\Xi_{P_{\bm{\alpha}}}$ [37, 38].
The value of $M_{n}$ (or $M_{1}$) can be estimated using Eq.s (11,12) averaging over a finite number of samples $\mathcal{S}$; the accuracy of the method of course increases with the number of samples. Indeed, the difference between the average value of $M_{n}$ obtained by sampling the distribution $\Xi_{P}$ and the actual value $M_{n}$ is of the order of $O(1/\sqrt{S})$. In addition, in the typical cases where the SREs is extensive, the number of samples required to estimate the SREs within a given error scales polynomially with $N$ for $n=1$, and it is exponential in $N$ for $n\neq 1$. [37, 38, 39].
It is worth mentioning two additional methodological aspects. Perfect Pauli sampling for MPS can be enhanced by employing a clever biasing strategy, enabling direct access to the stabilizer nullity [59], which can be seen as well as the $n\to\infty$ limits of the SREs [36]. Additionally, improved sampling strategies can also be applied to SRE, as demonstrated in Ref. [60].
@@ II.4.3 Replica Pauli-MPS
Figure 2: Phase diagram of the $S=1$ $XXZ$ chain with uniaxial single ion-type anisotropy. The three marked points correspond to the transitions examined in the rest of the paper: (i) Large D-XY ($J_{z}\sim-0.183$, $D\sim 0.5$); (ii) Haldane-Large D ($J_{z}\sim 0.5$, $D\sim 0.635$) and (iii) Haldane-Néel ($J_{z}\sim 2.93$, $D\sim 2.6$). $m_{1}$ is computed with perfect Pauli sampling with $N_{S}=10^{3}$ samples.
@@ III SRE in spin-1 XXZ chain
where $S^{\alpha}$’s, $\alpha=x,y,z$, are the spin-1 operators, $J_{z}$ is the easy-axis anisotropy, and $D$ is the single-ion anisotropy. The model has a global $U(1)$ symmetry corresponding to the conservation of total magnetization $\sum_{i}S^{z}_{i}$, and here we consider the ground states at the sector of zero total magnetization. The model displays a rich phase diagram, as sketched in Fig. 2, making it a good playground to explore magic in phase transitions [61].
@@
Fig. 2, obtained utilizing perfect Pauli sampling, illustrates the variation of $m_{1}$ within the model’s phase diagram. Notably, magic reaches its peak within the topological phases (Haldane and XY) and rapidly declines in the trivial phases (Ferromagnetic, Large D, and Néel). However, no distinct characteristics are observed at the critical point, consistently with previous findings [39].
@@ III.1 Scaling of full-state magic
[…]
@@
[…]
@@ III.2 Mutual information and long-range magic scaling
As concerns long-range magic, the numeric results are challenging to interpret due to statistical error. In Fig. 9a and Fig. 10a, we display the scaling with the system size $N$ for case $BC$ and $AC$, respectively. Comparing with the behavior of mutual information, long-range magic appears to increase less rapidly for two connected partitions, while remaining constant for disconnected partitions. Turning to the scaling with the bond dimension $\chi$, depicted in Fig. 9b, Fig. 10b, we notice parallels with the behavior of mutual information. However, unlike for the full-state magic case, we are not able to predict a functional form for the correction in the bond dimension. Specifically, to estimate the $\chi$-dependence, in the previous section we focused on $\chi\geq 5$. Yet, at these values of $\chi$, it is difficult to discern the scaling, as the fluctuations we aim to fit are of the same order as the error bars with a sample size of $N_{S}=10^{6}$. This is because our estimation of long-range magic comprises a sum of two terms, each associated with its own statistical error, and the $W$ term defined in Eq. (16) is more challenging to estimate. As a consequence, we cannot predict the precise scaling of long-range magic in $\chi$, but we consistently observe rapid saturation.
@@ Acknowledgments
We thank A. Hamma, T. Haug, G. Lami and L. Piroli for useful discussions and feedback on the manuscript, and M. C. Bañuls, R. Fazio, G. Fux and X. Turkeshi for collaboration on related topics. P.S.T. acknowledges support from the Simons Foundation through Award 284558FY19 to the ICTP. M. D. was partly supported by the MIUR Programme FARE (MEPH), by the EU-Flagship programme Pasquans2, by the PNRR MUR project PE0000023-NQSTI, and by the PRIN programme (project CoQuS). M.C. was partially supported by the PRIN 2022 (2022R35ZBF) - PE2 - “ManyQLowD”. MD would like to thank the Institut Henri Poincaré (UAR 839 CNRS-Sorbonne Université) for their support. MD and ET were partly supported by QUANTERA DYNAMITE PCI2022-132919,
Note added: While completing this work, we became aware of a related study in the context of random matrix product states [67] by Lami, Haug and De Nardis, that appeared on the same arxiv posting.
@@ Appendix A Autocorrelations
To further illustrate the origin of the above results, in Fig. 13, we show the behavior of the normalized autocorrelation function of the mutual information $I$ for case $BC$. Figs. 13a, 13b and 13c correspond, respectively, to the Large D-XY, Haldane-Large D and Haldane-Néel transitions. Counterintuitively, we observe that despite the mutual information increasing with size (see Fig. 7), the autocorrelation of the Pauli-Markov chain remains constant.